This page: http://monitor.mikes.fi/ionclock/

Notes - Comb - Ion - Cavity - DAQ

Archived notes: index_2019 index_2020 index_2021 index_2022spring
Machines and IPs:
Strontium/linux .45.99 (strontium.ad.vtt.fi) (192.168.10.1) - since 2021-12-03 IP = .45.15
win: .45.23 (ESPTE1M012-ioni.ad.vtt.fi)
kavi: .45.190 (kavi.ad.vtt.fi)
comb: .45.15 (192.168.10.18)
N210 SDR: (192.168.10.20)

News and notable results

Experiment/Programming TODO:

2023-01-03

Summary of Bartington measurements (sensor on optical table on ASE-side)
Keysight DMM measures Bartington voltage output at ca 2 S/s
peaks at 1e-3...1e-2 Hz are train rep-rate?


14:30local stop RF on/off loop, log-file entries are:

UTC 2022-12-25 08:51:58.807997 LOCAL 2022-12-25 10:51:58.807998 SET 10.0 REAL +1.0000000000000E+01
UTC 2022-12-25 08:58:04.703273 LOCAL 2022-12-25 10:58:04.703274 SET 10.0 REAL +1.0000000000000E+01
UTC 2022-12-25 22:58:05.051472 LOCAL 2022-12-26 00:58:05.051475 SET 10.0 REAL +1.0000000000000E+01
UTC 2022-12-26 12:58:05.191314 LOCAL 2022-12-26 14:58:05.191317 SET 10.0 REAL +1.0000000000000E+01
UTC 2022-12-27 02:58:05.351460 LOCAL 2022-12-27 04:58:05.351463 SET 10.0 REAL +1.0000000000000E+01
UTC 2022-12-27 16:58:05.487463 LOCAL 2022-12-27 18:58:05.487467 SET 10.0 REAL +1.0000000000000E+01
UTC 2022-12-28 06:58:05.631455 LOCAL 2022-12-28 08:58:05.631458 SET 10.0 REAL +1.0000000000000E+01
UTC 2022-12-28 20:58:05.775460 LOCAL 2022-12-28 22:58:05.775463 SET 10.0 REAL +1.0000000000000E+01
UTC 2022-12-29 10:58:05.959463 LOCAL 2022-12-29 12:58:05.959465 SET 10.0 REAL +1.0000000000000E+01
UTC 2022-12-30 00:58:06.099461 LOCAL 2022-12-30 02:58:06.099464 SET 10.0 REAL +1.0000000000000E+01
UTC 2022-12-30 14:58:06.243470 LOCAL 2022-12-30 16:58:06.243475 SET 10.0 REAL +1.0000000000000E+01
UTC 2022-12-31 04:58:06.407464 LOCAL 2022-12-31 06:58:06.407466 SET 10.0 REAL +1.0000000000000E+01
UTC 2022-12-31 18:58:06.583469 LOCAL 2022-12-31 20:58:06.583473 SET 10.0 REAL +1.0000000000000E+01
UTC 2023-01-01 08:58:06.759485 LOCAL 2023-01-01 10:58:06.759489 SET 10.0 REAL +1.0000000000000E+01
UTC 2023-01-01 22:58:06.902481 LOCAL 2023-01-02 00:58:06.902484 SET 10.0 REAL +1.0000000000000E+01
UTC 2023-01-02 12:58:07.243466 LOCAL 2023-01-02 14:58:07.243469 SET 10.0 REAL +1.0000000000000E+01
UTC 2023-01-03 02:58:07.466522 LOCAL 2023-01-03 04:58:07.466525 SET 10.0 REAL +1.0000000000000E+01

2023-01-02

10:47local adjust variacs, +2 DIVs on each
V2 to 44V
V4 to 56V
V6 to 56V
10:50local picture of Vescent thermistors
ch1 40.0915 C
ch2 39.7695
ch3 39.5827
ch4 39.5763
GP reads 3.3e-7 mBar

15:02local
ch1 44.3032 C
ch2 41.4037 
ch3 39.9249
ch4 39.8546
GP 3.8e-7

15:04 Change B-field measurement to Y-axis, along long edge of optical table.
(no change in datalogging, goes to same file as previous X-measurement)
reads ca +24 mV = 2.4 uT now

2022-12-30

12:25local stop Z-direction B-meter

turn B-sensor to measure along short edge of optical table, X-direction
12:32 restart measurement, reads ca +1.2 mV

2022-12-30_X_Bfield_1d.txt
2022-12-30_X_Bfield_1w.txt


2022-12-25

10:58 local: Start 14-hour trap RF cycling (while True)

2022-12-23

Stop coil-current and reference-voltage logging.

10:30local Continue with B-field measurement, now in Z-direction

2022-12-23_Z_Bfield_1d.txt
2022-12-23_Z_Bfield_1w.txt


2022-12-22

try to fix cavity temp-control and logging script (gpib timeout problem)
run every 6 minutes from crontab now (was every 5 minutes)

16:20local close cavity doors (but many temporary cables prevent closing doors properly)

2022-12-20

Measure I2 current directly and Vref from current driver:
data: 2022-12_I2_DMM_1d.txt (directly in A)


data: 2022-12_I_Vref_1d.txt

2022-12-19

11:33local stop B-field measurement

I2 = -11.4941 mA (box front-panel display)  ARTIQ software set-point is -11.4939 mA
thru 50 Ohm shunt-resistor should give ca 550 mV voltage?

data: 2022-12_I2_thru50R_1d.txt (divide measured voltage by 50 Ohm to get nominal set current -11.4941 mA)

2022-12-16

Measure B-field in direction of clockrun-Bfield with Bartington Mag-01H

Keysight DMM reads voltage into influx ca 2 S/s

voltage ca -44 mV
Bartington display -4.4 uT

Clockrun B-field direction
chamber coordinates b = [0.25228  -0.96341  -0.090579]

16:43local turn off clocklaser, comb (leave comb osc)
try opening closing blue lab-toolingcart drawers - seems to change B-meter step-wise?
16:52local leave lab
1 day and 1 week B-field voltage plots (multiply voltage by 100x to get uT)
2022-12_Bfield_1d.txt
2022-12_Bfield_1w.txt


2022-12-12

ca 17 local: first measurement of two clock-lasers against each other
old cavity locks comb, comb DDS adjusted down from 35 MHz to get new cavity beat-note a bit lower
beat-note from new cavity around 27.5 MHz
17:12 1S/s data collected about 15+ minutes (one interruption in the middle)

Strong 24? Hz oscillation, from scroll-pump?

collect 5 mins of 1000 S/s data also: "2022-12-12_sdrdata_5min_1000.txt"

2022-12-12_sdrdata_5min_1000.txt

2022-12-09

Second beat-note detector built.
Fermionics FD80FC, OPA818 op-amp, 5k1 transmipedance. Design: https://github.com/aewallin/One-Inch-Photodetector/tree/master/OPA818_ALTpinout
postgain ZFL-1000LN+
coupler for -10dB monitor output: ZEDC-10-2B
splitter for RF1 RF2 main outputs: Z99SC-62-S+


2022-11-28 measurements with old detector at 1348nm

2022-12-07

shaking test with feedforward correction?
150 mVpp from siglent siggen into audio PA (volume 1-2 clicks from max)
lower shaking amplitude chosen, since it gives sin()-shaped motion as seen by seismometer

f = 10...30 Hz in 1 Hz steps
sdr tracker data 1000 S/s for 30s at each shaking frequency
scope collects voltages, 100 ms/DIV, 30 kpts/trace
ch1: speaker drive, 50 mV/DIV
ch2: vtt seismometer Z voltage, 100 mV/div

run14: feedforward on, gain = -2560
(possibly wrong buffering in sdr-code...)

run16: repeat, ffw gain -2560
run17: ffw OFF

search for better gain, maybe -1060
run18: ffw gain -1060
run19: ffw gain -860  (seems to work for 25..28 Hz maybe?)


FFW on g=-2560 : too high gain, worse than with no ffw?
FFW off:
FFW on g=-1060 : Note gain optimized at 25Hz shaking, and this shaking suppressed quite well!?
FFW on g=-860 :


some sanity-check data with spekaer off:

2022-12-05

Continue with shaking-tests and feedforward?

moved speaker closer to center of cavity, above pumping-T

turn off 4 GHz scope (behind comb-rack)

log cavity-transmission voltage with DAQ, ca. 3.06 V now
shaking tests show similar results Saturday and Monday, below 30 Hz. In data from Monday there is a large step after 35 Hz?

2022-12-03

16:06 connect GP-pressure gauge cable (audio plug) to kavi measurement system

shaking-test run9
siglent siggen 200 mVpp into audio PA, volume on max
freqs 15...40 in 1 Hz steps
scope channels:
ch1 speaker drive
ch2 seismometer Z
ch3 cavity transmission
ch4 PDH detector dc-level (low-pass filter)
1 min of SDR tracker data per shaking frequency
500ms T/DIV for collected scope voltages, 30 000 pts/12 TDivs on screen

NOTE: previously shaking tests with big speaker had additional weight on top of speaker moving part (steel plate, fixed with M8 bolt)
- this additional weight was removed at some point when the speaker was not on top of the cavity
- velocity/acceleration amplitudes might be smaller with no added weight?

attach the steel plate as added weight
run11 is shaking run with added weight, 200 mVpp from siggen, 15...40 Hz in 1Hz steps
-> now clear peak in SDR tracker FFT, corresponding to laser-frequency fluctuation at speaker-drive frequency (at least above 20 Hz)

run12 repeat run, no autoscaling for seismometer voltage. check that seismometer cable to artiq does not cause attenuation if signal.
19:24 leave lab
slow recovery after repairs, Friday-Saturday:

"run8" shaking-loop test from ca 2022-11-24: (same 200mVpp speaker drive)

new result 2022-12-03, laser frequency measurement not very reliable, probably limited by intrinsic beat-note noise (from comb?)
Note <6e-11 sensitivity result for all freqs now!?

2022-12-02

Seems the Drift AOM was mounted loosely/wrong.
re-mount AOM, with spacers under 1" post to mount it slightly higher
re-align double-pass driftcomp AOM, and light into cavity
try locking to sidebands to evaluate sideband-to-carrier ratio -> raise EOM drive by +4 dB from -18 dBm to -14 dBm

after this rebuild, try virbation-tests and feedforward correction again?
- before lunch and AOM-repairs, feedforward seemed to work briefly (shaking cavity with speaker at 20 Hz)
- after AOM-repairs results not as clear, acceleration-sensitivity of cavity improved?

- cavity recovers from repairs very slowly during Friday-Saturday, with large negative drift settling towards zero

2022-12-01

Cavity output photodetector -3 dB bandwidth maybe only 12 kHz
rise/fall-time 0.35/BW = 29 us

measured ca 110 us (?)

FWHM = 1 / (pi*tau) = 2.9 kHz
Finesse = pi * (FSR) * tau = pi * (501 304 785)*tau = ca 173000

PDH LO angle 218.6 deg
lock to sidebands with 180 deg change from this


2022-11-29

Try modulation on BOA amplifier, to modulate intra-cavity power
measure laser frequency with 1000S/s tracker

"mod_run1"
modulation at 12 Hz
amplitudes 0.5 Vpp, 1.0, 2.0, 3.0, 4.0, 5.0 Vpp

PDH-detector shows larger variations, cavity-output detector smaller

leave running with 100 mHz modulation, 1 S/s tracker data should see this?
5Vpp from siglent, via divider-box to BOA mod-input, gives currents: 483 mA ... 525 mA
data from ca 11:55 local

mod_ampl    cav_amp_rel   pdh_amp_rel   pdh/mod  cav/mod  cav/pdh   laser(Hz)
0.49418     0.008         0.0135        0.02736  0.01622  0.59279   28.5
0.98682     0.0168        0.0267        0.02707  0.01702  0.62864   37.4
1.47672     0.0256        0.0399        0.02705  0.01733  0.64062   20.5
1.96933     0.0316        0.0538        0.02730  0.01604  0.58736   9.0
2.44901     0.0417        0.0663        0.02708  0.01703  0.62902   2.5

conclusions:
- ca 2.7% power-modulation per volt on PDH-detector
- ca 1.65% power-modulation per volt on cavity-transmitted detector
- ratio of detector modulations ca 0.6 output/pdh  - but note that detectors have different gain!
- no reasonable signal on laser frequency seen!? try larger modulation?

BOA modulation input-divider used R1=560 Ohm, R2 = 5.6 kOhm -> div-ratio R1/(R1+R2) = 0.09
remove divider to get more modulation?

no clear peak in frequency FFT?
leave with 2 Vpp amplitude, 20 mHz frequency
'slow_mod.txt' has voltage readings from this run
running until 16:30 
16:55 stop BOA modulation
'slow_mod_off.txt' voltages with modulation off

average cavity-photodiode ca 3.7 V ( previously 3.3V <-> 1.7 uW on ouput of cavity)
pdh-photodiode ca 1.3 V average ( previously laser-unlocked 1.6V and 25 uW)
mod_ampl  cav_v   cav_rel  pdh_v    pdh_rel  cav_rel/pdh_rel
2.44901   0.15473 0.0417   0.08511  0.0663   0.62902
laser frequency fluctuation 6.224 Hz

->
p_out_uw = 0.0797 uW (amplitude of power-fluctuation at cavity output)
p_pdh_uw = 1.33 uW (amplitude of power-fluctuation at cavity input)
y_laser = 6.224/(0.5*SRS) = 2.8e-14
p_out_uw / p_in_uw = 0.06

->
shift from cavity output-power
 shift/output-power = 2.8e-14 / 0.0797 = 3.5e-13 / uW = 160 Hz / uW
 shift/input-power = 2.8e-14 / 1.33 = 2.1e-14 / uW = 9.4 Hz / uW

TLi shutting down pumps from ca 17
connected new Gamma ion-pump to lab network, possibly 192.168.10.55


1000 S/s sdr tracker, pumps on
ON
OFF


2022-11-28

clocklaser + comb running during weekend, seems reasonably stable (but lab temperature oscillates)

measure beat-note SNR and detector noise
VLFX-80 lp-filter (DC-80MHz) normally on RF1-output removed

comb on: -21.9 dBm in fiber before detector (fiber detector)
laser (BOA 500 mA) + comb on: -5.6 dBm (tlabs big detector)
only comb again: -23 dBm (big detector)

2022-11-24

shaking run 7
collect cavity transmitted photodiode on ch3
PDH detector DC-level on ch4

shaking frequency 15...40 Hz in 1 Hz steps
200 mVpp from siglent siggen

run8:
no autoscaling of votlage ranges
cavity transmission and PDH DC-level seem strongly anti-correlated
some modulation seen on these at higher frequencies
shaking at 15...40 Hz, 1 Hz step
15 Hz, 500 ms timebase, 12 Divs, so 6s of data. 30kpts so 5 kpts/s
ch3: AC-coupled, transmitted power (usually 3..4VDC?)
ch4: AC-coupled, PDH dc-level (usually 1..1.5 VDC)

1.7 uW cavity transmitted power, with Thorlabs slim sensor, before big cube

PDH-detector DC-level:
1.6 V laser unlocked
1.3 V laser locked
24-25 uW on PDH-detector, with laser unlocked
-> ca 5uW into cavity
note 1.7uW cavity-output in both directions, so 3.4uW total out.

transmitted photodetector DC level ca 3.3 V

2022-11-23

Effect of fiber-noise on measured acceleration sensitivity?

"run4" data
measure on cavity->ion fiber, with fiber-noise servo locked. (actual fiber that affects measurement is laser->cavity, similar fiber on similar route in lab)
VCO 178.5MHz mixed down with DDS= 167.5 MHz -> resulting 11 MHz signal collected with SDR-tracker

modulation frequencies
[0, 20...60 in 1 Hz steps, 0]  0=no modulation applied (at start and end of run)
scope ch4 connected after 24Hz measurement (start of 25 Hz measurement)
scope 100ms/div for low frequencies, 30 kpts/waveform
200mVpp from siglent siggen, into audio PA with volume on max. 

second run, "run5" data
now collecting laser frequency as usual with SDR-tracker

null-test dummy run "run6" - same as run5, but audio PA off so no shaking applied

fiber noise
laser frequency
laser frequency, speaker OFF

2022-11-22

SDR tracker data with different sampling rates
2k, 1k, and 500S/s agree for tau<0.1s - seems ok
100 S/s fails spectacularly because strong 100 Hz fourier component is aliased down to 0 Hz!? (see also time-series with slow oscillation)
1 S/s data is surprisingly good?
5 minutes, ca 300 s

5min 1000S/s data
5min 2000S/s data

500S/s, note 85Hz peak (from oscilloscope on top of cavity-box??)
This peak probably an aliasing-effect - disappears at 1000S/S!?


beatnote 1000 S/s sample time-series and FFT spectrum
ca 100 kHz peak-peak on 1348nm - main component is at 100 Hz


Old IonClock project acc.sensitivities, from static tilting of the setup

Zoomed in view of run3 data (2022-11-21) shows ca 0.5e-10 [1 / (m/s**2)] around 22 Hz .. 25 Hz

2022-11-21

Shake test "run3"
200mVpp from siglent siggen, 20 Hz to 60 Hz in 1 Hz steps, audio PA on max volume, large speaker on top of cavity.
Now collecting X, Y, Z zesmometer channels also on scope
ch1: speaker drive (fixed V/div scale, always 200 mVpp from signal generator)
ch2: Z (autoscale)
ch3: Y (autoscale)
ch4: X (autoscale)
only one side of seismometer differential signal collected

scope autoscale voltage/div, collects 30kpts per waveform.
all triggered at once, so phase-relation should be ok?

1 minute of 1000 S/s sdr-tracker data, so 60 kpts collected at each shaking frequency

acceleration in m/s**2
modulation frequencies: f = [20., 21., 22., 23., 24., 25., 26., 27., 28., 29., 30., 31., 32.,
       33., 34., 35., 36., 37., 38., 39., 40., 41., 42., 43., 44., 45.,
       46., 47., 48., 49., 50., 51., 52., 53., 54., 55., 56., 57., 58.,
       59., 60.]
ax = [0.00060319, 0.00064918, 0.00083012, 0.00093529, 0.00091765,
       0.00108113, 0.00117883, 0.0013816 , 0.00156788, 0.00185419,
       0.00200961, 0.00203816, 0.00291352, 0.00304604, 0.00235276,
       0.00326701, 0.001061  , 0.00262235, 0.00686677, 0.01192614,
       0.01209723, 0.01074271, 0.01188085, 0.01815949, 0.03556015,
       0.09168198, 0.09170002, 0.05872096, 0.04274737, 0.03385898,
       0.03231065, 0.03121793, 0.03022256, 0.03015994, 0.02997894,
       0.03259281, 0.02823312, 0.03070945, 0.0331403 , 0.03783495,
       0.04262111]
ay = [0.00133874, 0.0016467 , 0.00209999, 0.00263785, 0.00315346,
       0.00371294, 0.00447876, 0.00517714, 0.00591638, 0.00672679,
       0.0076685 , 0.00904656, 0.01010779, 0.00929127, 0.00880376,
       0.00949138, 0.01035489, 0.01299954, 0.01625167, 0.01436025,
       0.01441011, 0.01499431, 0.0170222 , 0.01847578, 0.02415598,
       0.21742737, 0.02375181, 0.01247624, 0.0108871 , 0.01424358,
       0.01754852, 0.02069375, 0.02376994, 0.0254996 , 0.02979708,
       0.03376869, 0.03849049, 0.04606369, 0.05394652, 0.06505677,
       0.09408794]
az = [0.04208193, 0.04797343, 0.05446118, 0.06118383, 0.06831682,
       0.07562693, 0.08300574, 0.08993455, 0.09604655, 0.10097918,
       0.10508552, 0.10821076, 0.10997122, 0.10900603, 0.10714   ,
       0.10765576, 0.10889735, 0.10939354, 0.11028213, 0.10047205,
       0.0907674 , 0.08709093, 0.0828155 , 0.07760851, 0.06983236,
       0.05908808, 0.10236395, 0.26460852, 0.0888179 , 0.08412899,
       0.08199222, 0.08063443, 0.07935599, 0.07827483, 0.07706   ,
       0.07649192, 0.06786081, 0.06445975, 0.06251488, 0.0576693 ,
       0.0559294 ]
measured laser frequency oscillation amplitude, at 1348nm
f_ampl = [  1079.408851,    818.723792,   -482.989049,   -488.06891 ,
         -602.91406 ,    692.829355,  -1064.25438 ,   1563.591577,
        -1997.486954,  -2770.328896,  -3454.893118,  -4558.908448,
        -5563.664229,  -6318.976453,   6605.869224,  -8099.236535,
        -8870.7012  , -10368.164612,   9313.764894,  15463.886949,
       -14312.719534,  15001.81362 ,  16061.564614,  18588.079757,
        16935.776509,  14597.851283, -12921.079711, -14645.762528,
        11826.277249,   7600.756416,  -8701.945076,  -9738.099133,
        -9235.953525,  -5486.863949,   4540.778818,   2623.893065,
         2523.738117,  -7158.431644,   9072.185509,   7811.02465 ,
         6559.927516]
abs() fractional frequency amplotude
y = [4.85014377e-12, 3.67879891e-12, 2.17023079e-12, 2.19305630e-12,
       2.70909384e-12, 3.11311324e-12, 4.78204968e-12, 7.02573815e-12,
       8.97537471e-12, 1.24480112e-11, 1.55239864e-11, 2.04846953e-11,
       2.49993979e-11, 2.83932675e-11, 2.96823724e-11, 3.63925695e-11,
       3.98590174e-11, 4.65876197e-11, 4.18498503e-11, 6.94843987e-11,
       6.43118198e-11, 6.74081492e-11, 7.21699637e-11, 8.35224384e-11,
       7.60980891e-11, 6.55930118e-11, 5.80587182e-11, 6.58082930e-11,
       5.31394058e-11, 3.41527322e-11, 3.91007399e-11, 4.37565254e-11,
       4.15002178e-11, 2.46543087e-11, 2.04032328e-11, 1.17900262e-11,
       1.13399966e-11, 3.21652195e-11, 4.07643536e-11, 3.50975374e-11,
       2.94759410e-11]
Z acceleration sensitivity: kz = y/az
kz = [1.15254793e-10, 7.66840969e-11, 3.98491359e-11, 3.58437225e-11,
       3.96548572e-11, 4.11640823e-11, 5.76110744e-11, 7.81205693e-11,
       9.34481772e-11, 1.23273044e-10, 1.47727169e-10, 1.89303682e-10,
       2.27326721e-10, 2.60474291e-10, 2.77042851e-10, 3.38045719e-10,
       3.66023749e-10, 4.25871761e-10, 3.79479876e-10, 6.91579360e-10,
       7.08534358e-10, 7.73997352e-10, 8.71454797e-10, 1.07620200e-09,
       1.08972529e-09, 1.11008873e-09, 5.67179368e-10, 2.48700578e-10,
       5.98296122e-10, 4.05956778e-10, 4.76883553e-10, 5.42653130e-10,
       5.22962613e-10, 3.14971078e-10, 2.64770733e-10, 1.54134275e-10,
       1.67106699e-10, 4.98996956e-10, 6.52074422e-10, 6.08599991e-10,
       5.27020539e-10]




laser on, correct cavity now at 47.7 V piezo voltage (down from 62V last week)
1000 S/s tracker, with pumps running, ca 11:30 local

With 40 Hz 200 mVpp speaker shaking on.

2022-11-18

Cavity shaking test with VTT equipment
Siglent siggen ca 200mVpp into audio amplifier -> speaker on top of cavity
Rigol 4-ch scope collects driving waveform, and seismometer Z-channel

SDR tracker at 1000 S/s collects data into file

excitation frequencies ca 20Hz to 100Hz

frequencues (Hz), acc. sensitivities ( 1 / (m/s**2)):
[20., 22., 25., 28., 30., 33., 36., 39., 41., 44., 47., 49., 52.,
       55., 58., 60., 63., 66., 69., 71., 74., 77., 79., 82., 85., 88.,
       90., 93.]
[1.11454298e-10, 4.70609446e-11, 4.37762124e-11, 9.50034631e-11,
       1.43215661e-10, 2.11444179e-10, 5.06509628e-10, 6.47183783e-10,
       7.00980816e-10, 1.06368236e-09, 6.80575099e-10, 6.78055822e-10,
       6.61432245e-10, 5.53996509e-10, 9.44780161e-10, 3.92886489e-10,
       1.14245532e-09, 2.14449158e-10, 1.96252838e-10, 7.31444432e-11,
       1.65677968e-10, 1.81329515e-10, 5.36073733e-11, 7.47100880e-11,
       1.49642771e-11, 7.69647078e-11, 2.28410753e-11, 1.80890276e-10]



Note 2nd and 3rd harminic visible in both seismometer and laser frequency signals. Lasers shows subharmonic(?) around 65 and 80-90 Hz?

2022-11-17

fast SDR tracker at 1ms gate or 1000 S/s
large frequency fluctuations seen at 50Hz and 100Hz might come partly (or completely) from comb also??

50Hz peak at 63 dB
100Hz peak at 67 dB

try modulating non-inverting FALC-input with test-signal and check if this is visible with fast tracker?
1 mVpp 30 Hz -> not visible? (with 1000 Hz/V sensitivity, this should give 1Hz frequency-oscillations)
10 mVpp 30 Hz -> ca unclear peak? (10Hz frequency oscillation?)
20 mVpp 30 Hz -> ca unclear peak??
100 mVpp 30Hz -> clear peak in falc-output now, 
200 mVpp 30Hz -> no clear peak?

hard or impossible to create peaks in tracker spectrum by modulating non-inverting falc-input
laser-locking unstable with higher amplitudes than 200 mVpp -> 200Hz or smaller frequency modulations not seen in tracker-spectrum


2022-11-16

tests with fast sdr tracker and speaker shaking cavity

at 500S/s, see 50Hz and 100Hz peaks in fft of frequency time-series??
also 34-35 Hz from A/C, and many other peaks (pumps?)

2022-11-15

11:40 clock-laser on. now correct cavity mode at higher piezo-voltage of ca 62 V
beat-note 15.441 MHz

2022-11-14

ca 11:30 accidentally turned off turbo-pump.., pump front panel T=21C, ca >1300Hz
closed valve
11:44 now turbo starting again, with closed valve, "starting"
11:49 pump starting 3/6 bars, p=2.9e-8
11:53 4/6 bars, p 2.3e-8
11:58 5/6 bars, p 1.9e-8
12:00 "normal" 1350 Hz, p=1.9e-8
open valve
12:02 p up to 3e-6 with valve slowly opened
12:03 2.4e-6
14:32 8.8e-7



Lab temperature now in sustained oscillation mode?
2021-12-22 similar result -0.000548 V/C (average voltage 0.0586 V) -> -0.94%/C
Now -0.000342 V/C (average ca 0.05955) -> -0.57%/C

2022-11-11

09:50 relock laser to TEM00, now PDH dc-voltage ca 1.66 V

17 local: turn off clocklaser (comb remains on, Jaani uses)
PDH-discriminant proportional to square-root of optical power.
Power inclreased ca 1.5-fold, sqrt(1.5)=1.22 - measured increase ca +15%

2022-11-10

Try to measure PDH errosignal Hz/V with slightly modified settings.

Before changes PDH detector DC-output was ca 1.2 VDC

Increas BOA current to ca 600mA, now PDH detector ca 1.7..1.8 VDC
This increases overall loop-gain, so use FALC main-gain trimmer to reduce overall loop gain slightly

Detector signal is now 1.5 fold
expect error-signal scaling to go from 1061 Hz/V to 707 Hz/V!?

16:56: start 30 mHz modulation voltage run, 1h per amplitude setting
ampl=[0.05, 0.1, 0.15, 0.2, 0.25, 0.3]

Siglent Technologies,SDG2042X,SDG2XCAC2R0212,2.01.01.23R7
C1:BSWV WVTP,SINE,FRQ,0.03HZ,PERI,33.3333S,AMP,0.05V,AMPVRMS,0.017675Vrms,OFST,0V,HLEV,0.025V,LLEV,-0.025V,PHSE,0
C1:BSWV WVTP,SINE,FRQ,0.03HZ,PERI,33.3333S,AMP,0.05V,AMPVRMS,0.017675Vrms,OFST,0V,HLEV,0.025V,LLEV,-0.025V,PHSE,0
C1:OUTP ON,LOAD,HZ,PLRT,NOR
2022-11-10 14:56:12.069159 
2022-11-10 14:56:12.069159 C1:BSWV WVTP,SINE,FRQ,0.03HZ,PERI,33.3333S,AMP,0.05V,AMPVRMS,0.017675Vrms,OFST,0V,HLEV,0.025V,LLEV,-0.025V,PHSE,0
2022-11-10 15:56:12.223432 
2022-11-10 15:56:12.223432 C1:BSWV WVTP,SINE,FRQ,0.03HZ,PERI,33.3333S,AMP,0.1V,AMPVRMS,0.03535Vrms,OFST,0V,HLEV,0.05V,LLEV,-0.05V,PHSE,0
2022-11-10 16:56:12.379439 
2022-11-10 16:56:12.379439 C1:BSWV WVTP,SINE,FRQ,0.03HZ,PERI,33.3333S,AMP,0.15V,AMPVRMS,0.053025Vrms,OFST,0V,HLEV,0.075V,LLEV,-0.075V,PHSE,0
2022-11-10 17:56:12.531447 
2022-11-10 17:56:12.531447 C1:BSWV WVTP,SINE,FRQ,0.03HZ,PERI,33.3333S,AMP,0.2V,AMPVRMS,0.0707Vrms,OFST,0V,HLEV,0.1V,LLEV,-0.1V,PHSE,0
2022-11-10 18:56:12.683435 
2022-11-10 18:56:12.683435 C1:BSWV WVTP,SINE,FRQ,0.03HZ,PERI,33.3333S,AMP,0.25V,AMPVRMS,0.088375Vrms,OFST,0V,HLEV,0.125V,LLEV,-0.125V,PHSE,0
2022-11-10 19:56:12.839443 
2022-11-10 19:56:12.839443 C1:BSWV WVTP,SINE,FRQ,0.03HZ,PERI,33.3333S,AMP,0.3V,AMPVRMS,0.10605Vrms,OFST,0V,HLEV,0.15V,LLEV,-0.15V,PHSE,0

More power on the PDH detector increases pdh signal spectrum slightly:

2022-11-09

modulating non-inverting input of FALC
17:05local start siglent signal-generator, 30 mHz modulaiton, increasing amplitude once per hour
 0.05 Vpp, 0.1, 0.15, 0.20, 0.25, 0.30, 0.35, 0.40, 0.45, 0.50, 0.05 Vpp
11 hours total, ready 04 am in the morning?

Siglent Technologies,SDG2042X,SDG2XCAC2R0212,2.01.01.23R7
C1:BSWV WVTP,SINE,FRQ,0.03HZ,PERI,33.3333S,AMP,0.05V,AMPVRMS,0.017675Vrms,OFST,0V,HLEV,0.025V,LLEV,-0.025V,PHSE,0
C1:BSWV WVTP,SINE,FRQ,0.03HZ,PERI,33.3333S,AMP,0.05V,AMPVRMS,0.017675Vrms,OFST,0V,HLEV,0.025V,LLEV,-0.025V,PHSE,0
C1:OUTP ON,LOAD,HZ,PLRT,NOR
2022-11-09 15:06:06.479258 
2022-11-09 15:06:06.479258 C1:BSWV WVTP,SINE,FRQ,0.03HZ,PERI,33.3333S,AMP,0.05V,AMPVRMS,0.017675Vrms,OFST,0V,HLEV,0.025V,LLEV,-0.025V,PHSE,0
2022-11-09 16:06:06.631440 
2022-11-09 16:06:06.631440 C1:BSWV WVTP,SINE,FRQ,0.03HZ,PERI,33.3333S,AMP,0.1V,AMPVRMS,0.03535Vrms,OFST,0V,HLEV,0.05V,LLEV,-0.05V,PHSE,0
2022-11-09 17:06:06.787434 
2022-11-09 17:06:06.787434 C1:BSWV WVTP,SINE,FRQ,0.03HZ,PERI,33.3333S,AMP,0.15V,AMPVRMS,0.053025Vrms,OFST,0V,HLEV,0.075V,LLEV,-0.075V,PHSE,0
2022-11-09 18:06:06.939442 
2022-11-09 18:06:06.939442 C1:BSWV WVTP,SINE,FRQ,0.03HZ,PERI,33.3333S,AMP,0.2V,AMPVRMS,0.0707Vrms,OFST,0V,HLEV,0.1V,LLEV,-0.1V,PHSE,0
2022-11-09 19:06:07.091433 
2022-11-09 19:06:07.091433 C1:BSWV WVTP,SINE,FRQ,0.03HZ,PERI,33.3333S,AMP,0.25V,AMPVRMS,0.088375Vrms,OFST,0V,HLEV,0.125V,LLEV,-0.125V,PHSE,0
2022-11-09 20:06:07.243441 
2022-11-09 20:06:07.243441 C1:BSWV WVTP,SINE,FRQ,0.03HZ,PERI,33.3333S,AMP,0.3V,AMPVRMS,0.10605Vrms,OFST,0V,HLEV,0.15V,LLEV,-0.15V,PHSE,0
2022-11-09 21:06:07.396352 
2022-11-09 21:06:07.396352 C1:BSWV WVTP,SINE,FRQ,0.03HZ,PERI,33.3333S,AMP,0.35V,AMPVRMS,0.123725Vrms,OFST,0V,HLEV,0.175V,LLEV,-0.175V,PHSE,0
2022-11-09 22:06:07.551437 
2022-11-09 22:06:07.551437 C1:BSWV WVTP,SINE,FRQ,0.03HZ,PERI,33.3333S,AMP,0.4V,AMPVRMS,0.1414Vrms,OFST,0V,HLEV,0.2V,LLEV,-0.2V,PHSE,0
2022-11-09 23:06:07.703434 
2022-11-09 23:06:07.703434 C1:BSWV WVTP,SINE,FRQ,0.03HZ,PERI,33.3333S,AMP,0.45V,AMPVRMS,0.159075Vrms,OFST,0V,HLEV,0.225V,LLEV,-0.225V,PHSE,0
2022-11-10 00:06:07.855442 
2022-11-10 00:06:07.855442 C1:BSWV WVTP,SINE,FRQ,0.03HZ,PERI,33.3333S,AMP,0.5V,AMPVRMS,0.17675Vrms,OFST,0V,HLEV,0.25V,LLEV,-0.25V,PHSE,0
2022-11-10 01:06:08.007444 
2022-11-10 01:06:08.007444 C1:BSWV WVTP,SINE,FRQ,0.03HZ,PERI,33.3333S,AMP,0.05V,AMPVRMS,0.017675Vrms,OFST,0V,HLEV,0.025V,LLEV,-0.025V,PHSE,0


Note: comb measures IR, and we compare modulation voltage peak-peak to laser frequency peak-peak - a factor 2 was missing yesterday (amplitude vs. peak-peak)

Typical data with 350 mVpp (HiZ) = 350/2 mVpp (50R)
Measured frequency amplitude 46.774 Hz amplitude IR = 2*46.774 Hz amplitude red = 2*2*46.774 Hz peak-peak red
Sensitivity 2*2*46.774/ (0.35/2) Hz/V = 1069 Hz/V

Try the same measurement with some changed parameter? EOM-drive level, Toptical input level, Optical power?

2022-11-08


17:09local: 100 mVpp modulation on non-inverting falc-input, 10 mHz, ca 25 Hz oscillation in 1S/s comb data IR frequency?
-> PDH error signal slope ca 50 Hz (red) / 100 mV = 0.5 Hz/mV
17:37 amplitude 100 mVpp, 10 mHz
17:43 to 20 mHz, 300 mVpp
Note 200mVpp and 300mVpp rigol signal-generator settings, FALC 50 Ohm input sees half of this
Toptica EOM-drive +18.2 dBm, input level -10 dBm
PDH-detector ca 1 VDC


in-loop PDH errorsignal, when laser locked.
60 dB gain LNA, 50R resistor sets scale of voltage noise
LNA powered from Siglent +/-12V, replace with battery to reduce possible 50Hz signal from PSU?

2022-11-07

Trap RF on/off during weekend:

2022-11-02

Continue with helres measurement.
with probe -> 6 dB attenuation -> ZFL-500LN -> phasemeter (reads +5.2 dBm at 14.424 MHz)

5 min data trace:
- small pertubation by hitting optical table
- large
- try clockrun-program (ca 120s to 210s) (80ms probe, 35ms cooling, shutter shakes trap/breadboard slightly?)
- 

2022-11-01

Measure helres-probe with Microsemi phase-meter?
probe connected to SMA-antenna at side of helres

Probe -> 10 dB att -> Mini-Circ ZFL-500LN amp -> Phase-meter (measures +2.0 dBm @ 14.424 MHz)
note signal is quite sensitive to vibrations, hitting the optical table, or helres itself
This corresponds to the usual 3.44 Vpp measured by oscilloscope (10x probe to oscilloscope 1 MOhm input, 10x-mode)

Compare to direct 14.424 MHz output from Keysight siggen
normal setting of 10 Vpp is too large for phase-meter, reduce to 5 Vpp, phase-meter measures +18.1 dBm



Note roll-off in helres PN after 10 kHz - this is filtering effect from the Q of the helres+trap resonance?
Weaker effect in AM also

2022-10-31

Try to measure noise in coil-driver?
I2 = ca -11 mA, from driver directly to Keysight DMM measures -11.488 mA
(setting is -11.4939 mA in ARTIQ-program)

current through a 50 Ohm load-resistor, DMM measures -579.3 mV voltage over resistor

set I2 to -20 mA
measure voltage over 50 Ohm resistor with USB-soundcard
reference: 50 Ohm resistor with nothing connected
DMM reads -1.00822 V over resistor

Note with laptop-charger connected to 230VAC (back of shelf, above coolinglaser)
-> huge spikes at 50 Hz and harmonics, out to 3 kHz or so

test soundcard voltage input scale with signals from Rigol siggen:
2 mVpp (Hi-Z) 52 Hz
4 mVpp
8 mVpp
100 mVpp 


2022-10-26

RAM testing?

PDH EOM drive -18.2 dBm, phase 205.7 deg (with no bias-T before EOM-resonator)

Connect Bias-T before EOM-resonator, to create RAM with a slow DC-modulation?
ca +15 dBm power at 25 MHz measured after Mini-Circ PA + LPF
Toptica source -18.2 dBm, mini-circ ZX60-100VH+ (+36 dB typ., +33 dB min) -> +17.8 dBm typ (+14.8 dBm min)
(+30 dBm at 1dB compression, so 15 dB room to increase RF-drive)
LO angle adjusted to 218.6 deg after inserting Bias-T

Mini-Circ Bias-T is ZFBT-4R2GW (?) max DC voltage 30 V, but seems to survive much higher voltages.

Used Falco single-channel piezo-driver with gain of 20x and max output +/-175 V
could see small 1 mV or so modulation on PDH errorsignal (low-pass filtered) with laser unlocked
modulation from Rigol signal generator 1Hz or 0.1Hz with amplitude 1...15 Vpp (20x falco gain)

also hard to detect change in 25 MHz peak height with either PDH-detector or RAM-detector

2022-10-25

09:30 coolinglaser in noisy mode
signal ok with small adjustment to TPA current
also adjusting TPA temperature can change from noisy to normal mode.

ion lost?
TPA seems broken, output facet and/or metallization on top of TPA at output end damaged?


2022-10-24

844nm laser output ca 13 mW
<100nW in reverse direction of DLI-1 (so 50 dB isolation)
ca 11.5 mW in forward direction of DLI-1

P422 ca 3.5 mW, with round Si-detector after IR-splitter(for old fiber->old comb)


17:16 coolinglaser now OK again? (at least for a while)
      ion is still stored
      slider pos6

uMotion: A 3.2%, B 4.5%, C 10.9 %
RID25336 run uMotion-minimizer, 30 kphotons per measurement
print:V = [-6.4700 0.1943 22.3670 63.8193 ]
after optimization: A 0.9%, B 0.3%, C 7%

TPA went spontaneously into noisy mode!?
Can tune between normal/ok-mode and noisy mode with TPA-current (either up or down) to find normal working mode.

18:14 start 80ms clockrun
      all beams were on... -> switch to B -> p_exc improves
      all components p_exc ok, Dsum ok. 

18:33 switch to normal/bug interleaved run



After adjustments, reasonable Rb-signal found again:

2022-10-23

Fix possible bug in drift and f_cavity computation.
These need comb-data (1S/s) and drift AOM data (ca 1S/60s)
old code used numpy.interp() for comb-data that was outside the time-interval where drift-AOM data exists
e.g. numpy.interp(x_comb, x_dds, f_dds) uses last value of f_dds-array for any x_comb > x_dds

new code only computes fcavity for time-interval where drift-AOM data is available.
in use from ca 10:45 local
from 10 min of comb-data, 600 points, this causes 0-1 min (0-60 points) at the end to be excluded from drift/cavity computation

plot also residual drift on comb-page

2022-10-22 kiky Saturday

10:30 relock to Rb-line restores fluorescence. ion was dark from maybe 19:30 yesterday.

B-beam counts
pos1 301.9(15.5) / 3.6(1.8)
pos2 48.7(6.1) / 3.3(2.0)
pos3 82.2(9.8) / 3.6(2.0)
pos4 86.7(8.6) / 3.5(2.1) repeat 89/3.1
pos5 21.5 / 3.4
pos6 128.1 / 3.4 (use this for clockrun-test and interleaved run)

10:34 start 80ms normal clockrun, to find transitions
      pmt-threshold is 14 (old setting)
      all transitions excited, so B hasn't changed much from yday
11:39 end normal clockrun

11:39 start interleaved normal/bug clockrun
      coolinglaser intensity should gice ca 120 counts / 10ms
      80ms probe, 30 shots, g1=1.0, g2=0.01
      Probe DDS attenuation: artiq 18dB, RCDAT 25 dB
      Note there is still no 674nm power-stabilization for 674nm delivered to trap breadboard (todo: change to larger photodiode, and shield it from stray light)
      12 servo cycle-time 90s

13 ion dark. maybe Rb-unlock again..


Problems with Rb-signal (again)
After doubler, P422nm ca 1.2 mW
leave lasers on, PMT off, ion trapped?



Note with magcheck vertically close to camera, coolinglaser shutter during clockrun gives clear spikes in magcheck signal:
Polarity changes depending on shutter open/close
Horizontal scale is 50ms/DIV. 80ms probe, 45ms dead-time (35ms repumper, 30ms cooling, but some padding in clock-cycle also).

2022-10-21

Move Vbias/Icoil PSU to reduce AC B-field?
coils: I1 = -0.0408, I2 = -11.4939, I3 = -0.7741
latest Vbias V = [-6.4700 0.2351 22.5187 63.2567 ]

Magcheck on shelf above trap, reads 7.7 mVpp with PSU on

all beams on, slider to pos1 - try to keep ion trapped with zero bias-voltages?
decrease voltages gradually to zero
ion still stored with zero voltages, but image is smeared out, side-to-side movement in camera plane, a lot

Vbias/Icoil PSU off, magcheck now reads 0.97 mVpp  (reduction ca 7-8 fold)
with PSU placed on floor, under 461nm laser (where ITC laser-driver is now) - no apparent increase in magcheck reading when PSU switched on
power on Vbias/Icoil again
ramp back to nominal voltages

uMotion now: A 3%, B 2.3%, C 4.6%

back to nominal coil currents I1 = -0.0408, I2 = -11.4939, I3 = -0.7741

cavity jumped frequency -700 Hz, why? (ca 59873.57)

18:16 try to start normal clockrun? 80ms probe
18:42 restart with higher B, from guess based on C1- peak lock and comb-linecenter
      all transitions now excited, B-estimates agree, Dsum ok
      maybe moving psu had effect on static B-field also..

19:40 ion dark
      full power, all beams, 405nm?
      



08:55 interleaved clockrun still running, servo-cycle 515
      drop in p_exc, lowest point 08:40 - then higher again?
      servo-cycle length 91.7 s, slider movement with 91.7/2 s period causes one sample of invalid comb data with this period
09:35 stop interleaved clockrun
      leave pos2/OD1 filter

13:18 ion dark. all beams, max power, 405nm on
      relock to Rb-line -> fluorescence restored

No significant line-center frequency difference between normal/OD1 cooling intensity and high/OD0 seen


Excitation probability as moving average of 10 servo-cycles (10*30 = 300 shots in 10 servo-cycles)

09:38 uMotion after clockrun: 3.9/3.3/9.8 %

lab temperature now quite stable? try detuning scans
1000ms PMT counts
A 3232.8(54.1) / 252.5(13.6)
B 4626.6(282.6) / 248.6(13.2)  repeat 4696.8(432.8) / 278.4(16.7)
C 3560.3(27.3) / 248.8(16.1)
repeat B: (fit fails but data RID25199 ok?)

Magcheck sensor on lab floor, outer shield top (pump end), shelf above trap/below Vbias/Icoil PSU
top-of-shield: 2.043/0.111 = 18-fold larger 50Hz AC magnetic field compared to floor

2022-10-20

ion dark, stop clockrun
clearout on, 405nm on
slider pos1
detuning scans
07local no ion
all beams -> ion recovers!
slider pos2
using slider causes Rb-lock to jump to wrong line!?
back to pos1 to get usual fluorescence counts
continue normal/bug clockrun, now with 422nm locked to wrong Rb-line and 10-fold more power

10:33 local, stop clockrun
with slider pos1 and wrong Rb-line, fluorescence is 65.6 / 3.1
pos2: 9.5 / 3.2

pos1, correct Rb-line: 265.4 / 4.5 (back lab lights now on)
pos2 61.4 / 5.1

Elliptec slider was at max velocity '64' - this caused Rb-unlock when changing position
Try half speed '32' ?
no visible change in slider speed..

mount slider to optical table (not touching 422nm laser breadboard)
B-beam fluorescence counts:
pos1: OD0       313.8 in 10ms
pos2: OD1.0     49.3
pos3: OD0.9     75.2
pos4: OD0.8     87.1
pos5: OD1.3     19.4
pos6: OD0.7     124.8

15:25local, pos2, uMotion: A 6.3%, B 6.1%, C 5.2%
A-beam B-beam C-beam

Prepare interleaved clock-run with low and high cooling-laser power

Start with normal clock-run to find transitions?
16:30 normal clock-run started
16:48 restart with C2 B-value and linecenter
      C2 seems to lock ok - others not so good
17:11 restart again with C2 liencenter/B
17:41 C2 ok, also C1-. restart with C2 values, try 80ms probe
18:04 restart again, try old idea where FWHM probing-points are expanded initially during first 20 servo-cycles?
      C1, C4 one side locked to 50 Hz sideband -> linecenter shows 25 Hz offset
18:53 restart again with C2 values

19:26 restart again with larger B1 4.819177 uT
      finally all peaks seem locked and ok
      after ca 25 servo-cycles, peak servos have stabilized

19:48 start interleaved run with normal clock using OD1 filter, and comparison-clock using OD0 filter for 422nm
      Note 0.5s wait after changing slider position.
      OD1 ca 50-70 counts/10ms
      OD0 ca 290-310 counts/10ms

2022-10-19

NOTE: interleave-clockrun during the night (2022-10-18..19) does not show expected bug-behaviour where first shot of bug-clock would be always bright
Expected frequency-difference between interleaved clocks is zero.

(some restarts before this)
10:13 restart interleaved clockrun, now with fixed code that should reproduce the bug-behaviour!?

NOTE2: after AOM-chirp measurements 674-power stabilizing photodiode (on trap breadboard) was not aligned
so recent ion experiments are run without power-stabilization

from 14:30local p_exc seems to be dropping? why?
16local p_exc goes back up?
fairly large temperature-excursion in maser-room
this causes predicted line-center to diverge from measured ion line-center
Overnight interleaved clockrun (both interleaved clocks same, so expect no frequency difference)


From 10:13 restart, interleaved clockrun looks ok
left normal-clock, right bug-clock (no p_exc in first shot). 120 min of data:

With 4 hours of data:



Temperature-swing in maser room causes predicted line-center to diverge from measured line-center??

2022-10-18

Attempt interleaved run with clock-cycle-bug that leaves cooling/repumper on after n_shots probed.
This results in the next probe-pulse always giving a bright state (due to light-shift from cooling/repumper)

There are 12 servos, where we leave the cooling/repumper on as follows: ('E'-naming from previous EQS-runs)

0  C1+ normal-clock, NO cooling/repumper left on
1  C1- , NO
2  C4+ , NO
3  C4- , NO
4  C2+ , NO
5  C2- , YES leave cooling/repumper on
6  EC1+ bug-clock, YES
7  EC1- bug-clock, YES
8  EC4+ bug-clock, YES
9  EC4- bug-clock, YES
10 EC2+ bug-clock, YES
11 EC2- bug-clock, NO cooling/repumper left on

ca 16:43 local, start of interleaved run:
servo-cycle length doubles to ca 105 seconds
17:40 turn off lab-lights, PMT dark counts should improve
19:38 restart with slightly modified code (break realtime before leaving cool/repump on?)
19:51 restart again (remove default argument value)

08:30local ion still stored
uMotion A 1.9%, B 1.7%, C 4.2%

relock to Rb-line
now B-beam fluorescence better bright 172 / dark 4, in 10ms
B-beam 422nm ca 4 uW

slider pos2, OD1 for 422nm, B-beam ca 0.5 uW
bright 83.4 / dark 4.3 in 10ms

maybe fluorescence problem was locking to wrong Rb-line..
RID24966 is detuning scan with B-beam, 10ms gate
RID24967 is detuning scan with B-beam, 100ms gate
RID24968, B-beam 1000ms gate
B 1000ms gate counts: 7740.7 / 584.4

A-beam, improve alignment a bit
power 0.7uW
RID24971 A-beam detuning scan
A 1000ms gate counts: 4237.9 / 593.3

C, align
power 0.8 uW
RID24977 detuning scan with C-beam 1000ms gate
C 1000ms counts: 8024.9 / 581.3


B-beam
A-beam
C-beam

uMotion A 6.6%, B 5.6%, C 11.9%
uMotion optimization
V = [-6.4700 0.2351 22.5187 63.2567 ]

now uMotion: A 1.8%, B 1.9%, C 3.3%
14:25 restart clockrun with new B-estimate from B1
g1=1.0, g2=0.01, t_probe=100ms, artiq-att 18 dB, RCDAT 25 dB, n_shots=30
servo-cycle 52.24 s, dead-time per shot 45.1 ms
14:40 servos seem locked, D-sum ok, B1/B2/B4 consistent

Test continuous line-center computation, where (C1+C2+C4)/3 is computed after each pair is probed, based on last three pairs
-> new field 'linecenter' in clockrun database/file
add RID-number to database also (see when program has been restarted)
-> new field 'rid' in clockrun database/file


12:55 fixed ARTIQ issues, try 100ms clock-run

2022-10-17

15:30local: Prepare to load ion.
Set compensation voltages to (common-mode/diff) [-6.47, 0.26, 23.43, 59.5]
Check 461 dispenser. OK with -194.88 mA, 4.642 kOhm.
Optical powers: 
405nm: 4.8 mW
422nm: A 2.7 uW, B: 2.0 uW, C 2.4 uW
461nm: 33 uW 
clearout: 48 uW (lower than previously?)
repumper: 9 mW (slim-sensor with attenuator..)
trap RF-monitor 3.4Vpp (osc-probe at side of helres)

12:58:15 try to load with 1.75 A
         PMT background ca 15 counts/10ms
         422nm detuning 208 MHz

set detuning 197 MHz
empty trap with RF off, on
13:18 ion! out-of-focus and not aligned with PMT (bottom left corner of camera image)
align camera image to get PMT-counts (top left corner as usual)
image is reasonably sharp, 200 counts/10ms with all beams and full power
to 208 MHz detuning
B-beam align to ion (was quite ok, slight improvement in fluorescence)
check ASE-alignmnent, was ok

uMotion A 15%, B 34%, C 27%
Try Nelder-Mead uMotion minimizer? 20 kphotons, 10 iterations
V = [-6.4700 0.3083 22.7181 62.3532 ]
uMotion  A 0.8%, B 1.0%, C 4.9%
photon-counts seem quite low, compared to 422nm power in beams (B-beam up to 4 uW?)
TODO: Try focusing ASE-beam (low fluorescence common to ABC-beams)

Try Nelder-Mead again, 40 kphotons, 10 iterations
V = [-6.4700 0.2482 22.5815 62.9131 ]
A 1.6%, B 1.8%, C 0.6%
B bright 57.1, dark 4.0, threshold 15 (10ms gate)

Update AHM3 model from AB2022-10
Initial B-estimate 4.82 uT, line-center estimate from comb + AHM3-model

ARTIQ-attenuation for Zeeman AOM fixed at 18 dB (use RCDAT to change power)

RID24789, peak-search with 300us probe (Fourier limit 1333 Hz AOM), RCDAT 0 dB
C1+ Center =  75.77598233720198, height 38, FWHM 1363 Hz AOM
C1- Center =  75.74893607778606, height 49, FWHM 1312 Hz
B1 estimate =  4.819323980398982
C2+ Center =  75.78938645227757, height 53, FWHM 1396
C2- Center =  75.73549972012783, height 38, FWHM 1297
B2 estimate =  4.816239762220009
C4+ Center =  75.82981567098327, height 52, FWHM 1356
C4- Center =  75.69506914992346, height 37, FWHM 1273
B4 estimate =  4.814242777524356
C4 Linecenter =  75.762442
comb estimate:   75.762420
RID24789, 300us peak-search


ca. 5-6 dB of RF-attenuation with RCDAT should be equivalent to OD1 attenuation for optical power?
RID24790 rabiflop on C1+, seems 300us was very close to tau_pi
set RCDAT to 5 dB, this should give OD1 optical attenuation?
RID24792 rabiflop on C1+, visually tau_pi ca 950 us
set RCDAT to 10 dB, should give OD2 optical attenuation and tau_pi 3 ms?
RID24794 rabiflop on C1+, tau_pi ca 2.6 ms

RID24796 is peak-search with 10 dB RCDAT and 2.6 ms probe (Fourier 154 Hz AOM), 30 Hz step and 1 kHz span, 33 pts/peak
C1+ Center =  75.77593749649549, height 32, FWHM 186
C1- Center =  75.74891067377892, height 52, FWHM 155
B1 estimate =  4.8158605901426705, C1 Linecenter =  75.762424
C2+ Center =  75.78936094291716, height 49, FWHM 155
C2- Center =  75.73548414135105, height 31, FWHM 196
B2 estimate =  4.815352195469657, C2 Linecenter =  75.762423
C4+ Center =  75.82980227736321, height 48, FWHM 143
C4- Center =  75.69505262970313, height 39, FWHM 158
B4 estimate =  4.814354485141597, C4 Linecenter =  75.762427
RID24796, 2.6ms probe, 10 dB RCDAT (With ARTIQ set to 18 dB, and RCDAT to 10 dB, this is roughly equivalent to the old max-attenuation setting used in ROCIT-campaign)


RID24797 is rabiflop on C1+ with 10 dB RCDAT, tau_pi ca 2.6ms as before
set RCDAT to 20 dB - we now expect ca 30 ms tau_pi?

RID24800 is peak-search with 20 dB RCDAT and 30 ms probe, 3 Hz AOM step
C1+ fit fails, but central peak ca 4-5 steps wide, so seems Fourier limited

RID24800, 20 dB RCDAT, 30ms probe (Rabi-sidebands close to 50Hz sidebands!?)
Abort scan to adjust ASE focus?

B-beam fluorsecence before adjustment 90-110 counts / 10ms
focus few mm 'out'
hard to find ion after re-focusing objective... counts down to 600/100ms now..

set RCDAT to 25 dB - this should give maybe 90ms probe?
RID24826, 25 dB RCDAT, 90ms probe, 1Hz step, 200Hz span, 201 pts/peak
19:38local leave lab


Continue with 100ms clock-run?
some ARTIQ issue prevents clock-run...


Note effect of drilling/roadwork outside MIKES, ca 11:55 to 13:50 local
beat-note jumps of ~60 kHz

2022-10-15


Sat ca 08:30 restarted dritcomp again (ran ok Friday afternoon/evening a few hours)

From Friday afternoon collecting 10 S/s tracker data with VTT tracker.

Extra breadboard mounted on top of cavity. Speaker mounted on top of breadboard.
PTB seismometer (side of cavity, at DC-voltage terminals), PTB tiltmeter on top of cavity (close to PDH EOM)

Input-side door maybe still open? check..

ca 14:10 local, close input-side door

Sunday:
10:50 changed to 1s tracker (from 0.1s)
some laser-lock issue again? why?

laser jumping out-of-lock related to high ULI/fine1 voltage from FALC?
try to reduce it with piezo-control via computer.
0.001 V piezo step whenever fine1 > 0.2 V ?

seems to go away when lowering the EOM drive-level by 2 dB to -18.2 dBm?
-could this be overload/clipping of the EOM RF PA?
-or overload of the Toptica PDH input-stage?


Intermittent laser-lock problem, PDH-signal and PDH-error completely disappear, and integrator pulls in one direction
Why always two blank periods of similar length?

2022-10-10

Fiber-noise detector and signal?

Detector + 2 pcs Mini-Circ amplifiers
ca 47 dB SNR with SA RBW 1 kHz
Signal -36 dBm, noise -82 dBm

reference arm ca 70 uW
at fiber-noise detector ca 50 uW
-- from reference-arm: 52 uW
-- from fiber/ion-arm: ca 0.3 uW

Both arms blocked: 178.5MHz signal -53.7 dBm !? (this is electrical feedthru??)
signal disappers when disconnecting AOM-drive VCO cable
alufoil around detector: -58 dBm (not much improvement?
(notes above on 178.5MHz, not actual detector signal at 357MHz...)


At 357 MHz: (RBW 1 kHz)
peak: -4.5 dBm, noise: -59 dBm, SNR: 54 dB
fiber/ion beam blocked: -26...-30 dBm
reference and fiber blocked: -56 dBm (very close to noise floor at ca -59 dBm)
only reference blocked: ca -56 dBm
both unblocked: -4.1 dBm
ion/fiber blocked: -27...-30 dBm
mirror after AOM blocked: -36..-46 dBm
input of AOM blocked: signal disappears
Wide SA view of fibernoise-signal. Note SHP-300 high-pass filter.
Narrow SA view of fibernoise-signal

Compressor service
file: 2022-10-10_FFT/seism_XY.txt
span 0 to 150 Hz, 1601 points
spectra collected with 10s interval

11:20 started FFT-analyzer measurements of Z and Y seismometer signals.
Seismometer is on Minus-K, minus-K is floating
one compressor-motor was probably on (atlas copco compressor has 4pcs (?) identical motor+compressor units inside it)
11:25 agreed with service-guy to keep all motors off for 15 minutes, while we gather measurement data
11:30 changed various FFT-settings, spectra should be ok now?
      peak at 42 Hz, both Z and Y
      Y peak at 66-68 Hz (intermittent high/narrow peak?)
10:45 (one or more compressor motors may be on, depending on service)

Note 75 Hz most probably from oscilloscope on top of cavity-box

14:09 turn off oscilloscope on top of cavity box. this makes 75 Hz peak disappear
      seems 2/4 units running on atlas copco compressor currently
try averaging on FFT-analyser?
14:13 N=5 averaging on
15:45 oscilloscope on again, to test clock-laser



2022-10-06

cavity doors open changes EOM-temperature and PDH LO-angle?
Was at 205.7deg

2022-10-04

Changed PSU from Thorlabs LDS12B to side-of-cavity banana-plugs for the 357 MHz fiber-noise detector
1 kHz and 500 Hz signals (old PSU) are probably artefacts from the (helper) ARTIQ 178.5 MHz + 500 Hz DDS-signal interfering
Possibly a small change in the 50Hz and 100Hz signals when servo is off? (but a new peak at 22 Hz or so??)

Fiber-noise measurements:
- freerun1k, AOM-signal from ARTIQ, with 500 Hz offset giving 1 kHz beat-note
- freerun2, repeat
- vco, AOM-signal from VCO, servo not locked (drifts quite a lot)
- locked, servo locked
- servooff, servo electronics power off, noise floor

Servo/VCO locked: level is far below noise-floor when servo is off? 
500Hz peak remains - why?
VCO unlocked shows 1 kHz broad spectrum with side-bands at 1.5-2 kHz

noise-floor with servo off shows small 50Hz and 100Hz peaks!?
noise-floor large peak at 1 kHz is the same as closed-loop 500 Hz peak??



PDH detector signal with spectrum analyzer   
- yellow: SA floor
- green: detector dark
- blue: detector bright
- red: locked to cavity
measurements through 10dB coupler

data: CSV file with same name as image


2022-10-03

Measure PDH-signals and fiber-noise signals with Sound-Blaster x-fi

PDH-signals:
- no (large) 50 Hz component in instrument noise-floor or actual signals
- PDH error-signal similar when laser unlocked and laser off (not much evidence for shot-noise-limit here?)
- clear 34 Hz peak from A/C
- audio-digitizer not very useful <10 Hz ?

Fiber-noise:
- reference 357 MHz as usual
- VCO replaced with ARTIQ DDS at 178.5 MHz + 546 Hz
- after double-pass this gives 1092 Hz beat-note on error-signal
- spectrum plotted with -1092 Hz frequency offset


2022-09-30

Add grounding-wire between FALC and DLCpro, this seems to reduce the 10ms spikes a lot (completely gone?)

Measure with FFT: (all traces 1601 pts, frequency 0 Hz to 120 Hz)
ch1: FALC monitor when laser locked
ch2: sesmometer Y, X, Z
FALC signals also with FALC-off, and with FALC-on but no input signal connected

Errorsignal (FALC-input) via BNC-T to FFT-analyzer
- laser locked
- laser unlocked
Measurement of 50 Hz peak with FFT-analyzer seems hard/impossible, since peak remains when FALC off and/or input-signal disconnected
34 Hz signal seems 'real', from KSK A/C-machine
Not much evidence for 100Hz modulation

2022-09-29

Measure FALC-Monitor signal with Keysight black scope, Mini-Circ LPF filter (11MHz?) cuts high frequencies

Try FALC on/off states, and FALC-ULI -> DCLPro Fine-in-1 cable connected/disconnected

Connecting the Fine-in-1 cable gives spikes with 10ms period in FALC-monitor. Not good!??

Measurements with no photodetector-(error)signal connected to input of FALC.
FALC power on.

FALC power off.

For comparison, scope-scale increased 500-fold to 500 mV/DIV, FALC-monitor when laser is locked:


Toptica win-program settings:
PDH freq 25 MHz, Amplitude -18.2 dBm
Input level -10 dBm, Lock level 0V, Phase 199.3deg
Piezo ARC: Fine-in-1, -0.5 V/V

2022-09-27

Try also KSK105 effect? N=5 averaging mode, so takes some time for spectra to change.
15:54 KSK105 on as normal
15:58 KSK105 off
      16:00 no 34 Hz peak in Y-signal
      16:01 reduced 34 Hz peak in Z-signal also. 30 Hz peak remains
40-41 and 43-44 Hz signals remain - some motor?
16:14 KSK105 back on

17:12 stop data collection
Clear reduction of peak at 34 Hz (Z stronger, Y weaker) as well as 57.5 Hz (Z,Y equal)




Spectra when turning off devices (oscilloscope off already): Siglent Spectrum-analyzer, Siglent PSU, black PSU, Thorlabs PSU, Oscilloscope on top of fft-analyzer
>5 dB reduction in 50Hz Y-peak - otherwise no major changes


Oscilloscope on top of cavity-box causes 75Hz peak of height 20-30dB

14:13 start on new file: 20220927.txt, 10s interval for spectra, no averaging
      75 Hz peak possibly from Rigol oscilloscope on top of cavity-box? (oscilloscope now off)
14:15  try oscilloscope on
       clear peak at 77Hz initially, then 76 Hz, then ca 75 Hz - fan in oscilloscope?
14:18 oscilloscope off
      14:18:20 spectra: 75 Hz peak disappears
14:21 spectrum-analyzer (Siglent) off, also on top of box
14:23 EOM RF-amplifier + orange fan off (on floor at input-end of cavity, RF-amplifiers needs fan-cooling?)
14:26 black PSU ("Aalto") off, powering cavity through-transmission power servo  
14:29 Siglent PSU off, powering fiber-noise servo, 674nm trap-breadborad-power servo
14:33 small rigol-scope on top of FFT-analyzer off
14:40 TK203 (kallion välitila, tuloilmakone) off (no VFD, so runs at 50 Hz?) - no major change in spectrum?
      Atlas Copco compressor shows 88 Hz acoustic peak with spectroid
14:44 TK203 back on
14:47 Thorlabs-PSU LDS12B (on floor) off, powering fiber-noise photodetector(?)
      Similar model PSU causes bad B-field near ion (was used for RF rectifier, at some point)
      Y-signal 50Hz peak drops to <-90 dBm, slight improvement?
      strong remaining Z-peak ca 43-44 Hz!? (compressor??) and 34 Hz (KSK105), and 57.5 Hz (??)
14:51 turn on N=5 averaging on FFT



Note no averaging during bottom half of image. Top half with N=5 averaging in FFT-analyzer.
Calibration waveform applied to FFT-input around 12:00


Trillium compact voltage output sensitivity is ca 750 V * (s/m)
measurement is velocity, need to differentiate to get acceleration?

Old IonClock project cavity acceleration sensitivities estimated:
  0.15 e-10 s^2/m  longitudinal
  0.13 e-10 s^2/m  transverse
  0.24 e-10 s^2/m  vertical

From ion-clock spectra, estimated 50/100Hz modulation
 Beta_50Hz = 0.9
 Beta_100Hz = 0.55
FM modulation depth (?) 0.9*50Hz = 45 Hz ? -> 1e-13 fractional

1e-13 / 0.24e-10 (s^2/m) vertical = 4e-3 (m/s^2) vertical acceleration amplitude a_ampl at 50Hz
a_ampl = vel_ampl*2*pi*f_mod
-> vel_ampl = a_ampl / [2*pi*f_mod] = 1.3e-5 (m/s) 

Test of FFT-analyzer scale.
Signal from Keysight signal-generator (trap drive), set to 45 Hz.

Keysight:    Oscilloscope:     FFT-ch2-peak:
10 mVpp      ca 20 mVpp        -30 dBm
5  mVpp      ca 11 mVpp        -36 dBm
1 mVpp       3.4 mVpp(noisy)   -50 dBm

dBm = 10*log10( [ Vampl/sqrt(2) ]**2 / Z / 1 mW )
Vampl = Vpp/2
Note Keysight Vpp is into 50 Ohms, but we use FFT-analyzer with 1 MOhm input impedance.
Vampl = sqrt( 10**[ dBm/10 ]  * 1 mW * 50 Ohm ) * sqr(2)

Peaks in Z-spectra                    Vampl         Vel_ampl=Vampl/ [750 (V*s/m)]
75 Hz   -75.8 dBm                     5.1e-5 V      6.8e-8  m/s
44 Hz   -80.3 dBm                     3.1e-5 V      4.07e-8 m/s
Peaks in Y-spectra                    Vampl         vel_ampl
50, 75, and 100Hz all at ca -82 dBm   2.5e-5 V      3.35e-8 m/s


Seismometer spectra from night, on minus-k, doors closed


2022-09-26

Effect of wedges?
- Z higher noise around 60 Hz, lower noise around 32-35 Hz
- Y slightly higher noise around 70 Hz, slightly lower around 30-35 Hz

Minus-K peak that was around 2.4 Hz (+ weak 2nd harmonic) last week, now around 10Hz (note also 2nd harmonic 22Hz) ??
from cables? or some details related to Minus-K floating?
remaining high peaks around 32Hz, 50Hz 

16:22 seismometer power was of?, turn on.
16:33 doors closed
16:34 spectra with doors closed
      Try KSK105 off? ca 16:40-16:42
      ch1 autorange on?
16:47 spectra ok?

start new file: 20220926_2.txt
Note Minus-K trace in these plots probably wrong (Seismometer had no power?)

New file: fft_test_10s_20220926.txt
Seismometer on stone from start of file 09:45 or so.
10:19 turn off seismometer
10:21 FFT-analyzer electrical noise-floor with Seismometer off, but cables connected.
10:28 seismometer on concrete, FFT-analyzed Z/ch1 input range changed 
10:29 first spectra on concrete, OK
10:33 ch1/Z input range back to sensitive setting (was overrange before)
10:35 ch1/Z to -14 dBm input range, ch2/Y is -4 dBm

13:56 seismometer off
      move onto stone
14:00 ch1/Z shows overvoltage as transients settle?
14:06 ch1/Z range to -6 dBm, averaging off
      highest Y-peak now ca 72 Hz, broad peak around 70 Hz 
14:10 N=5 averaging on
14:12 averaged spectra ok? (now ch1/Z settled, 'half' light off at -6 dBm setting)
14:14 to -12 dBm rage ch1/Z ('half' light on)
14:15 averaged spectra ok
14:18 remove wedges?
14:46 most wedges removed (one of 8 wedges broke, some of it remains (middle, towards back of lab)
14:47 ch1/Z range now -18 dBm (half-light on steadily)
15:00 seismometer off
15:02 on minus-K, floating
      ch2/Y to -14 dBm range

         
Effect of wedges seems not very clear/good? Need to repeat no-wedges measurements?


Seismometer on stone, from Saturday evening to Monday morning
Autocal-off on FFT-analyser seems to help keep labview data-collection running
file: fft_test_10s_20220924.txt (332 Mb)

2022-09-24 kiky-Saturday

fft file: fft_test_10s_20220924.txt
had to reset fft-analyzer, some settings probably changed...
0 - 120 Hz Frequency, 10s record length, 1 MOhm input impedance
FFT is average of 5
data collection interval now 20s

17:56 Seismometer on concrete
18:00 fix X-axis sot that frequencies start at 0 Hz
18:03 turn off seismometer
18:05 reset measurement - Seimsometer now on stone
18:07 first FFT collected, on stone
18:10 move seismometer to lab floor
18:11 reset FFT
18:12 first FFT collected, on floor

install wedges under stone!?!

18:16 Z-signal was overrange, chane input to avoid overrange
18:21 seismometer off
18:27 first FFT, on stone, with wedges

auto-cal off on FFT-analyzer (this causes timeout in labview-program?)

18:33 leave lab

2022-09-23

8mm Hex-key fits between stone and concrete, 8.5mm metal-block does not fit
34 Hz peak identified as coming from KSK105 (set-point 35 Hz)

2022-09-22

Datasheet: Trillium Compact
new fft file: fft_test_10s_20220922.txt

13:44 electrical noise-floor of FFT-analyzer, with nothing connected to BNC-inputs
      no major peaks at 50Hz or 75Hz remain
13:47 50R termination on both input channels
      13:48 initial transient after conencting 50R-terminators dies down. Z-spectrum slightly better comperade to nothing connected to ch1. Y-spectrum(ch2) similar.
13:50 back to seismometer signals.
13:53 open cavity door (large transients)
      Y-signal recovers quickly back to similar spectrum as seen before
      Z-signal takes long time to settle? FFT-analyzer overvoltage/oversignal?
14:03 set Z/ch1 to AC-coupled, now similar FFT-spectrum as before
      no averaging on FFT-analuzer
-------------
14:09 seismometer on stone, next to ion-pump controller
      sharp spikes in Y-signal?? 2Hz repetition rate, two spikes close together?
      14:14 spkes disappear??
      time-series shows large oscillations, especially Y-signal, at ca 50 Hz
      spectrum shows highest peak at 43-44 Hz

seems like stone has a resonance around 40-50Hz? (oscilloscope image from hitting stone with hex-driver)

14:26 FFT signals connected again. Y-spectrum high (>-60 dB) peak at 43 Hz (stone resonance??)
      ca 44 Hz and 2nd-harmonic 88 Hz also highe peaks in Z-spectrum
--------------
14:34 seismomenter down to concrete pillar
      Z-signal now has 2 Hz spikes (internal mechanism of seismometer, after it is moved?)
      Y-spectrum settles faster, seems ok 14:36, only 50Hz spike remains
14:42 signals look strange, so reboot seismometer (DC supply off, on)
14:45 averaging-mode, N=5, on FFT-analyzer
-----------------
16:30 data from 14:45 -> ca 16:30 should be fairly quiet, from concrete pillar
      ca 68Hz peak visible now also?
16:42 seismometer on lab floor, input-side of cavity
      peaks 40, 45, 49 Hz
16:48 reboot seismometer (Y-signal was at +10V?)
      16:49 spectra look settled
      Z-peak at 24-25 Hz, also 34 Hz (from KSK105?) 
17:00 back to concrete pillar
      17:04 -> FFT data looks ok?
      
Note 68Hz extra signal appears in afternoon ca 15 local (similar to yesterday)


Seismometer on stone (next to ion-pump controller). Response when hitting stone at other end, with hex-key handle
Note response has frequency in 40-50 Hz range!?


Motor-testing:

11:25 -1floor PK304A off (this motor caused problems 2022/02, damping towards ceiling and horizontal tubes improved 2022/02)
      spectroid on phone shows vibrations around 30-35 Hz from this motor? (VFD set-point is 35Hz)
ca11:32  PK304 back on
11:40  -1floor PK102 off (set-point probably 50 Hz)
ca11:47 PK102 back on

Spectroid shows strong 42Hz signal close to AtlasCopco compressor (-2floor)

from spectrum looks like one problematic motor set-point could be 44 Hz? (motor vibrates with some 'slip' from set frequency, maybe -1 Hz lower)
also one peak lower at 41-42 Hz?


LabView program file-writing a bit unstable, data collected until ca 19.
time-axis is not completely linear (program stops sometimes), but time-stamps should be accurate

17:20 -3floor KSK105 off
17:30 -3floor KSK105 on, 35 Hz
17:47 scroll pump on
18:02 turbo soft start
18:20 turbo full speed

image color scaling min=-130 dB, max=-75 dB

Average spectrum for all data, and spectrum with KSK105 on/off (ca 23 spectra, 10s/spectrum, averaged for each setting)

2022-09-21

Seimsmometer installed at side of cavity
Minus-K Z-direction resonance frequency adjusted to maybe 0.8 Hz, should be floating now.
Seimsmometer Z-signal to ch1 and Y-signal to ch2 of FFT-analyser
FFT-analyser set to 0-120 Hz frequency and 5x-averaging (check?)
Traces read to file over GBIP every 10 seconds
Intermittent strong signal around 68Hz - source unknown?
From ca 17:20 local to 17:30 local, KSK105 in -3floor was off (normal setting is 35 Hz lately)
 this is seen as lower noise around minute 80 in the plot.

2022-09-14

SDR-tracker ch2 LO-mixing frequency
Set to 215 MHz
CH2 frequency in MU:  60517119992791
artiq CH2 actual frequency in Hz: 214999999.999999851  (this is probably wrong? check that FTW is even!)

In ZMQ-receiver code (takes gnuradio output and writes to InfluxDB) we have:
ch2_mix = 214999999.999996304512   # from even AD9912 tuning word: 60517119992791-1

2022-09-08

Result of first counter-test with pi, lambda, omega counting.
Some issues with numerical offset in pi-counting still?

2022-09-07

Test SDR-software with outputs for pi, lambda, omega counting.
Chronograf dashboard "SDR counter weights"
ca 17:20 measuring output with no FM

17:35 start FM-modulation loop
FM-depth 50 kHz
modulation frequencies linspace(0.01,3,55)
1000 s sleep at each modulation frequency.

2022-09-06

Frequency-response test of SDR-tracker+counter, using 1 s gate-time
Measurement as described in Dawkins2007, which gives abs( W(f) ) from measured ADEVs with FM modulated input-signal
Note Vernotte2016 (17) is square of measured response.

Weight-function (for freqeuncy vs. time data) for Pi, Lambda, Omega counters, and corresponding frequency-response

2022-09-05

Try to characterize beat-note SDR-tracker.

Signal from Keysight (trap-drive) generator
f0 = 14.024 MHz
fmod = adjustable 0.1 Hz to 3 Hz ?
FM-depth = 50 kHz (quite large?)
Amplitude 700 mVpp (seems OK for N-radio?)

fmod = linspace(0,3,55)
lowest possible modulation frequency seems to be 1 uHz
start loop over modulation-frequencies at ca 17:30 local
1000s per modulation frequency
signal connected to SDR-channel 1

modulation frequencies are linspace(0,3,55):
[0.         0.05555556 0.11111111 0.16666667 0.22222222 0.27777778
 0.33333333 0.38888889 0.44444444 0.5        0.55555556 0.61111111
 0.66666667 0.72222222 0.77777778 0.83333333 0.88888889 0.94444444
 1.         1.05555556 1.11111111 1.16666667 1.22222222 1.27777778
 1.33333333 1.38888889 1.44444444 1.5        1.55555556 1.61111111
 1.66666667 1.72222222 1.77777778 1.83333333 1.88888889 1.94444444
 2.         2.05555556 2.11111111 2.16666667 2.22222222 2.27777778
 2.33333333 2.38888889 2.44444444 2.5        2.55555556 2.61111111
 2.66666667 2.72222222 2.77777778 2.83333333 2.88888889 2.94444444
 3.        ]

ca 17:45:30 change to 2nd modulation freq fmod 0.055 Hz

2022-09-06 ca 9:25local: turn off FM modulation
datafiles:
/59827_comb.txt
/59828_comb.txt


AOM-chirp weekend run
- when switching to a new attenuation, the RCDAT is a bit slow, so we end up with a bit of the old attenuation during the first repeat
- 0dB first repeat was with warm AOM, second repeat has gain adjustments, 3rd+4th should be ok.
-  2...18 dB seems ok 

2022-04-28: first AOM chirp measurement. Limited SNR due to 60 MHz detector, ARTIQ attenuation level 18...28 dB
2022-08-12: programmable attenuator RCDAT installed, larger attenuation range. Better SNR with old PDH-detector (modified for higher bandwidth)
2022-09-05: optical setup shielded from aircurrents with cardboard + plastic bag shield

2022-09-02

Back to AOM-chirp measurements.
ARTIQ DDS source at 76 MHz and 15 dB attenuation.
RCDAT adjustable attenuation from 0 dB up

B-SDR tuned to 151999576.99998077750205993652 Hz
400 Hz software-shift
Resulting beat-note at 23 Hz, recorded at 200 S/s

Friday 15 to Monday 07, ca 64 hours.
Each run is 30min on, 30min off.
Attenuations from 0 dB down to 28 dB
2 dB steps give 15 different attenuations [0:2:28]
-> 4 repeats per attenuation gives 60 hour measurement time

ca 14:56 local: start AOM-chirp run
15:57 lower SDR gain to 14dB (spectrum was strange looking before this) (start of 2nd repeat with 0 dB)


CEO DDS measrement at setting 35 MHz - 1 Hz
Predicted (nHz)   Measured (nHz)
-223.517          -228.4 (32.0)
Updated DDS quantization figure:


2022-09-01

CEO DDS measurement at 35 MHz + 1 Hz setting:
Predicted (nHz)    Measured (nHz)
+223.517           +222.1 (32.4) 
Results so far:

2022-08-31

Leave same measurement running with DDS-setting of 35.000001 MHz over night
Expected fact-fset(35 MHz + 1 Hz) is +223.517 nHz
(For an offset of 35 MHz - 1 Hz, the expected fact-fset is -223.517 nHz. this could be measured also to validate DDS model)


Result from 16h CEO_DDS measurement with Microsemi 3120A:
35 MHz output mixed down with 42 MHz LO from ARTIQ -> meausurement around 7 MHz (low-pass filter before phase-meter)

DDS-model with CLK=14*10 MHz and 48 bits predicts zero frequency-error for 35 MHz output

f_LO = 42000000.000001586973667  (as output from artiq, check for numerical accuracy?)

f_measured = 7000000.000001602 Hz (stdev 7.762e-6 /sqrt(N=54900) = 3.313e-8 Hz)

Final DDS-error measured is: -13.7 (33.1) nHz  - compatible with prediction of zero.

DDS quantization measurements so far:

2022-08-30

Yet another attempt at measuring the 35 MHz f0/ceo DDS
Mix down with ARTIQ 42 MHz (as below)
IF amplified with mini-circ amplifier (9 dB attenuator to avoid compression - some 2nd harmonic still present)
11 MHz LPF on IF-signal to avoid 35 & 42 MHz at input of phase-meter
Meausrement at 7 MHz with Microsemi 3120A phase-meter (+2.5 dBm signal as reported by Microsemi)
NOTE: ceo-SYNCRO DDS has setting of "14-mult"! this may mean 14*10 MHz DDS clock!?
ca 16:25local start 16h run
NOTE2: mult=14 giving 14*10 MHz DDS clock seems to give exact output at 35 MHz!
CLK = 140 MHz
bits = 48
DDS_out = CLK * FTW/(2**bits)
gives exact FTW corresponding to 35 MHz, since 35/140 = 1/4


Try measuring DDS-frequencies with N-SDR.

ch1: rep-rate DDS directly at 19.999854 MHz
ch2: mixed-down ceo DDS at 35.000000 MHz

Mixing LO from ARTIQ, 42 MHz nominal
  in MU: 11821949021848
  actual frequency in Hz: 42000000.000001586973667

-> ch2 measurement nominal 42-35 MHz = 7 MHz
NOTE: set mixing-frequency to zero in zmq-receiver (writes to influxdb), needs to be reset to old value before clock-run

valid data from ca 11:25 local into database
long cables to SDR, may require stable lab-temperature to get sane results?

16:10 local stop measurement. N-SDR measurement has significant bias at ~1 uHz level - seems not suitable for measuring 0.1 uHz DDS offsets currently.

2022-08-29

Try verifying f0/ceo DDS frequency using HP 53230A counter.
see gist read_counter.py on github?
Measurement has lots of noise and averages down very slowly - try N-SDR measurement instead?

20 h measurement of rep-rate DDS frequency over weekend.
DDS setting 19 999 584 Hz.

Frequency offset from nominal / nHz
Predicted   16h-run         20h-run
-171.363    -173.7 (5.9)    -165.4 (3.0)

weighted mean -167.105 (2.67) nHz


2022-08-26

around 16local: leave 60 hour run with campaign setting of 584 Hz. Should finish by Monday morning.

Summary of DDS quantization results
Actual DDS-frequencies are spaced 426 nHz apart. Figure shows quantization error for integer Hz settings around 19.999584 MHz

3120A phase-meter measurement of rep-rate DDS-output (back of comb-rack)
Reference is 10 MHz output from the menlo 'PLO'-unit, behind the screen


('584' is the setting used in the ROCIT campaign)
DDS-setting (Hz)    Predicted offset (nHz)    Measured offset (nHz)    Menlo-screen-indication   Predicted Comb-shift@1348nm

19 999 576           +18.626
19 999 577          +100.583
19 999 578          +182.539
19 999 579          -160.187
19 999 580           -78.231
19 999 581            +7.451
19 999 582           +89.407
19 999 583          +171.363
19 999 584          -171.363                  -173.7 (5.9)             "-1 Hz"                   -1.71e-16
19 999 585           -89.407                  -83.0 (35.1)             "-1 Hz"                   -8.94e-17
19 999 586            -7.451                  -8.5 (25.6)              "-1 Hz"                   -7.45e-18
19 999 587           +78.231                  +73.2 (13.7)             "0 Hz"                    +7.82e-17
19 999 588          +160.187
19 999 589          -186.265
19 999 590          -100.583
19 999 591           -18.626

DDS-model:
clk = 12*10 MHz  (10 MHz input, 12-fold clock-multiplier)
bits = 48 (48-bit tuning word)
f_dds = clk*FTW/2**bits
Recipe:
1. compute exact FTW based on desired setting: FTW_exact = f_set*(2**bits)/clk
2. round to integer binary numer: FTW_int = round(FTW_exact)   (or perhaps floor() or ceil()?)
3. compute actual output f_act = clk*FTW_int/2**bits
4. compute DDS error: f_act - f_set

The spacing between actual DDS-frequencies is clk*1/2**bits = 426.326 nHz
if we always round the exact ftw to the nearest integer the maximum error is 0.5*426.326 nHz = 213.163 nHz

Comb-lock: 1 GHz (4th harmonic) rep-rate compared to maser-referenced 980 MHz + f_DDS
-> rep-rate error is DDS-error/4
for 1348nm we measure against comb-mode n_comb = 889558
f0 = -35 MHz  (from similar DDS as rep-rate!)
f_IR = f0 + ncomb*frep - fbeat(ch1)             (**)

Evaluate fractional shift from DDS-error (rep-rate):
f_IR = f0 + ncomb*(245 MHz + DDS/4) - fbeat(ch1)
DDS = f_set + f_err
this gives an f_IR error of:
f_IR_err = ncomb*f_err/4
fractionally
y_IR_err = ( ncomb*f_err/4 ) / (0.5*SRS)
The prefactors together are 1e-9 [1/Hz], so:
y_IR_err = 1e-9 [1/Hz] * f_err [Hz]    (compare to numbers above, -171 nHz DDS-shift gives -1.71e-16 fractional comb-shift)

For f0/ceo a similar DDS is used.
The model predicts a +141.561 nHz error for the 35 MHz DDS (set frequency of 35.000000 MHz) (so far we did not verify this with measurement)
This contributes with opposite sign, because Eqn (**) has -35 MHz 
The shift at 1348nm is y_IR_err_ceo = -141.561 nHz / (0.5*SRS) = -6.37e-22  (negligible compared to other errors)

In conclusion it seems the ROCIT campaign used a DDS-setting where the comb-mode(s) used are shifted -1.71e-16 from nominal.
A correction of +171 (e-18) should be applied to data in post-processing.
From a 16 hour phase-meter measurement this number is verified as +174(6) e-18.

2022-08-25

Lenovo laptop, USB3 to SDR. IP xx.xx.45.170
new artiq dds frequencies
ch1  MU: 70368718281966
ch2  MU: 70368718563441

2022-08-24

Sanity-check of SDR measurement
Replace comb rep-rate input with ARTIQ-generated DDS signal.
ch1 nominal 249.999942 MHz
actual: 249999941.999998748302460 (MU: 70368727852115)
ch2 nominal 249.9999 MHz (no change from past runs)
actual: 249999899.999998807907104 (MU: 70368716030166)

SDR tune 249999509.99678924679756164551 (no change)
SDR file 2022-08-24_check-1.dat

run2 with new SDR tune
SDR 249999499.99809250235557556152
SDR file 2022-08-24_check-2.dat

run3 with bandpass filters around measured frequencies.
ch1 bandpass around 42 Hz  (37-47 Hz, 5 Hz transition)
ch2 low-pass around 0 Hz (5 Hz cutoff, 5 Hz transition)
file: 2022-08-24_check-3.dat

run4, frequencies closer together?
CH1 frequency in MU: 70368717437541
CH1 actual frequency in Hz: 249999904.999999225139618 (artiq does not round to even FTW!!)
CH2 frequency in MU: 70368716030166 (no change)
CH2 actual frequency in Hz: 249999899.999998807907104
SDR tune: 249999494.99874386191368103027
ch1 SDR-note at +10 Hz
ch2 SDR-note at +5 Hz
file: 2022-08-24_check-4.dat

Reconfigure sampling and filters on SDR:
250 kS/s Hardware sample-rate
10 kS/s 1st-stage decimation (25x). Low-pass filter 4 kHz (80% of 5 kHz Nyquist-freq, transition 500 Hz, 10% of Nyquist)
1 kS/s  2nd-stage decimation (10x), Low-pass filter 500 Hz (transition 50 Hz)
200 S/s 3rd-stage decimation (5x), Low-pass filter 80 Hz (transition 10 Hz)
file: 2022-08-24_check-5.dat


Result from overnight run, with DDS setting of ROCIT campaign
SDR-tuning is different, giving larger 800 nHz offset for ARTIQ-reference
Final result is similar to previous value ca +200 nHz



2022-08-23

Change settings to ROCIT-campaign values for overnight run
DDS 19 999 584 Hz
SDR tune nominal: 249999510 Hz
SDR tune actual: 249999509.99678924679756164551 Hz
start of run 16:41 local

Comb DDS-setting 19999587 Hz.
Rep-rate measured from M-NIR branch with usual beat-note photodetector. Measures ca +220 nHz high from nominal
ARTIQ reference frequency measures ca +13 nHz high (note swings both below and above 0 due to temperature-changes)
Difference frequency ca +210 nHz high - corresponding to 8e-16 shift at SRS/2 (comb mode-number ncomb=889558)
Measurement from Monday afternoon to Tuesday. Image shows residual from linear fit to phase vs. time measurement.
Note significant slopes at start of measurement, from lab-temperature changes(?).



[Nemitz2021] notes on comb rep-rate measurement.
fr = 250 MHz comb rep-rate
fHM = 100 MHz maser reference

Maser multiplied by m=92 (Nexyn Corporation NXPLOS)
Comb harmonic n = 37 (DSC40S photodetector, Discovery Semiconductors)

n*fr = 9 250 MHz
m*fHM = 9 200 MHz
These mixed together, beat-note at 50 MHz to counter.
Counter accuracy requirement relaxed by factor n*fr/50MHz = 185

2022-08-22


Dual channel rep-rate measurement using Artiq DDS as reference, with settings:

CH3 set frequency MHz: 249.9999
print:CH3 frequency in MU: 70368716030166
print:CH3 actual frequency in Hz: 249999899.999998807907104
print:CH3: delta f, 1.1920928955078125e-06

Comb DDS at 19 999 587 Hz
Comb rep-rate 250+(20e6-DDS)/4 = 249 999 896.75 Hz

measure +2.749 uHz offset from expected rep-rate? (ca 6000s dataset, after throwing away first 3000s of data)
2nd run: +2.879 uHz

2022-08-19

Try comb rep-rate DDS setting of 19 999 587 Hz. This shows zero error on Menlo RRE-syncro??
17local start phase-meter run of length 24h
restart SDR-run, 2022-08-18-comb-reprate-200Sps-2.dat
SDR-tune  249999499.99809250235557556152 + software-shift +400 Hz

Monday 2022-08-22 stop SDR data collection.
analysis gives 8.9856e-8 Hz offset


Some offset seen between expected and measured comb rep-rate?
4th-harmonic of rep-rate is compared to phase-locked 980 MHz reference.
Resulting 20 MHz IF is phase-locked with programmable DDS.
This gives actual rep-rate as:

frep = 245 MHz + fDDS/4 = 250 MHz + (fDDS - 20 MHz)/4


Reprate-DDS       Phase-meter            SDR
                  ??? check divider??
19 999 584 Hz     +278(63) nHz          +278.17 nHz

2022-08-18

Test RF-lock of comb to maser.
Rep-rate set to clockrun-setting of: 250 MHz - (20MHz - 19 999 584 Hz)/4 = 250 MHz - 104.0 Hz
Rep-rate measured with usual beat-note photodetector (clock-laser is off) on the M-NIR comb-branch.

250 MHz beat-note measured with two instruments:

1)
250 MHz signal to Hittite HMC394 5-bit programmable divider, set to 25-fold division.
Divider output amplified with mini-circ amplifier to get suitable level for phase-meter.
Measurement at 10MHz with Microsemi 3120A phase-meter.
Expected result is (250 MHz - 104 Hz)/25 = 10 MHz - 4.16 Hz

2)
directly at 250MHz with B210 SDR
SDR-tuning at 249999499.99809250235557556152
software shift of +400 Hz
Samples recorded at 200 Sps

2022-08-17

Verify electrical RF-power delivered to AOM during chirp tests.
Signal chain: ARTIQ 15 dB, AA PA, RCDAT adjustable attenuator -> AOM
RCDAT   SA-measurement at 76 MHz
60      -31.5 dBm
50      -21.6
40      -11.9
30      -1.8
20      +8.2 dBm

fit to this data gives RF at AOM: +27.96 dBm - RCDAT-attenuation

Set new SDR-tuning frequency: 151999029.99972695112228393555

2022-08-16

10 dB RCDAT attenuation data collected during night
15 dB data during day.
15:45local For the last (?) repetition of 10, install plastic bag and cardboard shield to reduce airflow over beam-paths
run continues with 20 dB and 25 dB repeats during evening/night
16:25local start new SDR-file for the 20 dB run

2022-08-15

15:50local start of 10dB attenuation aom-chirp run (note first repeat differs from others)

2022-08-12

16:00local start chirp run, (1800s on, 1800s off) * 10 repeats * 6 attenuations [0, 5, 10, 15, 20, 25].
SDR datafile ca 6 Mb/hour * 60 hours = 360 Mb

Friday 4pm to Monday 9am is 65 hours. Time to do 6 different AOM attenuations, 10 hours each
Working attenation range is ca 0 dB to 25 dB, in 10 steps
ARTIQ attenuation set to 15 dB
mini-circ att,  SA peak height (optical beat-note at 152MHz, via splitter to SA)
0 dB    -13
3       -16.6
6       -21.1
9       -26.6
12      -32
15      -38
18      -43.5
21      -50
24      -55 (SNR only ca 10 dB now)

2022 Spring season archive: index_2022spring

New AOM-chirp measurement setup, now with mini-circ programmable attenuator after the AOM PA.
Old PDH-detector was modified with 3k9 transimpedance to give bandwidth suitable for measuring double-pass beat note at 150 MHz
Setup is similar to 2022-04 (2022-04-28) chirp-measurement, with added adjustable mirror to align beam onto detector
PDH-detector is followed by mini-circ amplifier (ca +23 dB) and splitter to SDR and Siglent-SA.
SNR for beat-note with maximal AOM-drive is >65 dB - clear improvement from old setup with 60MHz detector
142 Hz sideband @ -51 dBc
1.04 kHz sideband @ -44 dBc (this could be mechanical resonance of some component - sideband seems to jump up when hitting on the trap-breadboard?)

2022-08-08

Note 0.4 nW floor of power-meter restricts measurable range to ca 0-20 dB attenuation.
Measured power seems to agree with model, using optimal RF-power P0 = 1.3*Pmax_used
single-pass model is I/I0 = sin**2 [ (pi/2) * sqrt(P/P0) ]

small-angle approximation sin(x) ~ x for small x.
-> I/I0 ~ P single-pass, and I/I0 ~ P**2 double-pass

When this scaling holds, we expect 10 dB of reduction in optical power (OD1-filter) for each 5 dB of reduced RF power.


With RCDAT inserted after PA, 0dB setting, now ARTIQ attenuation ca 15 dB seems OK.

RCDAT  P674
0      3.8 uW
3      1.78
6      605 uW
9      177 nW
10     112
11     71.8
12     49
13     31
14     19.8 nW
15     12.9
16     9.1
17     6.0
18     4.0
19     2.6
20     1.9
21     1.3
22     1.0
23     0.8
24     0.6
30     0.4
40     0.4
50     0.4
off    0.4  (artiq DDS off)
Test 674nm power vs. AOM RF drive power, old configuration (no adjustable attenuator)
674nm power-servo set-point to 2.0 V, lab lights dark, detector-area covered with holy-blanket
Thorlabs slim detector after 461-cube, before B-fiber

Detector dark: 0.2 nW
18 dB DDS att   4.03 uW
19  3.45 uW
20  2.75
21  2.15
22  1.58
23  1.13
24  0.77
25  0.53
26  0.36
27  0.24
28  0.16
29 105 nW
30 67 nW
DDS off 0.5 nW
18 4.00 uW
17 4.35
16 4.42
15 4.24
14 4.08

AA Optoelectronics whitepapre on diffraction efficiency:
I1/I0 = sin**2 [ (pi/2) * sqrt(P/P0) ]
this formula for single-pass, we use double pass.
Note RF PA may go into saturation at high power-levels!



Earlier measurement of the RCDAT programmable attenuator: RCDAT-3000-63W2
Measured over typical AOM frequency range of 60-90 MHz.

Lowest dataset is SA floor

2022-08-03

Migration to ARTIQ-7 done.

2022-08-02

Starting migration to ARTIQ-7

Summer break 2022


For 2022 Spring results see index_2022spring

For 2021 results see index_2021

For 2020 results see index_2020

For 2019 results see index_2019